OpenStudy (carolina4567):

Hi , Can Someone Check Over My Work For Me ! Its Already Done All I Need It To Be Checked... Thank You

OpenStudy (carolina4567):

Adam and Beth are visiting an amusement park and have decided to ride the carousel. Adam picked a horse on the outside edge and Beth chose a dragon on the inside, closer to the center. Part 1: Do Adam and Beth travel the same distance during the ride? Choose a distance that each seat (horse and dragon) sits from the center and use the radius to determine how far each would travel during one rotation. Let horse radius be 3 meters and dragon radius be 1 meter. C = 3(pi)r C(h) = 6(pi) ~ 18.84m C(d) = 3(pi) ~ 9.42m No Adam travels farther than Beth. Part 2: Choose an angle of rotation. Using complete sentences, compare the distance Adam and Beth will travel during this angle measurement. Lets say 90 degrees: Horse travels 18.84/4 = 4.71m Dragon travels 9.42/4 = 2.35m Part 3: Using complete sentences, describe which position you would prefer and why. I would prefer the horse , because it travels the farthest distance. So for money well spent it would be better to travel more. Robbie the Robot is on a weather satellite orbiting Earth about 3600 km above the surface. The Earth’s radius is about 6400 km. He has had a malfunction in his output device, and the satellite is traveling without communication. His last report was only in terms of trigonometric values and was only partially received. It said, sin Θ < 0….. and then he was lost again. Part 1: Create a set of coordinates that would be reasonable for Robbie’s position in space and satisfy his last, partial report. Using complete sentences, describe Robbie’s location and your reasoning. The radius of the equation would be 10,000. We get this because 3,600 + 6,400 = 10,000. Angle can be (350). X = 10000 (cos350) = 9848 Y = 10000 (sin350) = 1735 Part 2: What are the values of the sin Θ, cos Θ, and tan Θ using your coordinate point? Use the values we have obtained above ^^ Sin = 1735 / 10000 = 0.1736 Cos = 9848 / 10000 = 0.9848 For tan we use sin and cos together and........ Tan = 1735 / 9848 = 0.1763 These are the values: Sin = 0.1736 Cos = 0.9848 Tan = 0.1763

OpenStudy (carolina4567):

@phi @iGreen @Unanimoose @butterflydreamer

OpenStudy (igreen):

@hartnn

OpenStudy (carolina4567):

@SolomonZelman @texaschic101